OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y =1/4 x + 2. y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (anonymous):

y=1/4x+2 m=1/4 m of new equation = negative reciprocal = -4 y=-4x + b point (-1, -2) -2 = (-4)(-1) +b -6 = b y = -4x - 6

OpenStudy (anonymous):

Thanks:D